Which of the following values represent the definite integral "\displaystyle{\int_{0}^{\pi/4}e^{\tan{\theta}}\sec^2{\theta}d\theta}"

1) "e ^2"

2) e - 1

3) e/2

4) e


1
Expert's answer
2021-08-25T08:59:15-0400

"\displaystyle \text{Let $u = \tan\theta$, when $\theta = 0,\, u = 0$ and when $\theta = \frac\pi4, \, u = 1$ and } du = \sec^2\theta d \theta
\\ \implies \int_{0}^{\frac \pi4} e^{\tan \theta} \sec^2 \theta \, d\theta = \int^1_0 e^u \, du = \left[e^u\right]^1_0
= e^1-e^0
\\=e-1"